# thanks: equal contributionthanks: equal contribution

### v1

@@ III Cellular appendages
In a letter sent to the Royal Society of London in 1676, van Leeuwenhoek reported the discovery of protozoa and the observation of cilia and flagella. He wrote: “I also discovered a second sort of animalcules, whose figure was an oval, … provided with diverse incredibly thin little feet, or little legs [cilia], which were moved very nimbly …, and wherewith they brought off incredibly quick motions” (cf. translation of letter 18 Dobell (1958); Haimo and Rosenbaum (1981)).
@@
Today, our language has changed a bit, and we describe cilia and flagella as slender, hair-like appendages that extend from the main cell body and perform diverse functions, such as enabling cell movement Beeby et al. (2020); Bondoc-Naumovitz et al. (2023), generating flows Eichele et al. (2020); Huang and Choma (2015); Ling et al. (2024), sensing environmental stimuli Marshall and Nonaka (2006); R. Ferreira et al. (2019) and attachment to surfaces Ryu and Matsudaira (2010); Jarrell et al. (2013). While there are a myriad of different filamentous appendages found in biology, here our focus is on those most widely studied, while also aiming to capture the diversity of their structures, properties, mechanisms and functionalities.
@@
Across the tree of life, the most well studied cellular appendages are those used for motility, in particular the flagella in prokaryotes and cilia in eukaryotes. These filaments whip, wave, or rotate through the action of molecular motors that transform the cell’s chemical energy into mechanical energy Beeby et al. (2020); Gilpin et al. (2020). The resulting non-reciprocal motion enables net propulsion within the cell’s microscale, viscosity-dominated environment Purcell (1977); Lauga and Powers (2009).
Box III | Drag anisotropy Drag anisotropy refers to the directional dependence of drag forces acting on an object as it moves through a medium. Due to their elongated form, filaments experience less hydrodynamic drag when pulled along their length than when pulled perpendicular to it. Drag anisotropy can be quantified as the ratio $\xi=\frac{\zeta_{\perp}}{\zeta_{\parallel}},$ (3) where $\zeta_{\perp}$ and $\zeta_{\parallel}$ are the friction coefficients for motion perpendicular and parallel to the object’s length. Typically, for filaments $\zeta_{\perp}\approx 2\zeta_{\parallel}$. The total drag force density (drag per unit length) opposing a filament’s motion at low Reynolds number (i.e. viscosity-dominated regime) is given by $\textbf{F}_{\text{d}}=-(\zeta_{\parallel}\textbf{v}_{\parallel}+\zeta_{\perp}\textbf{v}_{\perp})$, where $\textbf{v}_{\parallel}$ and $\textbf{v}_{\perp}$ are the projections of the local velocity into directions parallel and perpendicular to the segment Lauga and Powers (2009); Hancock (1953).
@@
A low Reynolds number world. At the microscale in which these filaments operate, inertia is insignificant and viscous forces dominate. In fluid dynamics, this is the world of ‘Low Reynolds number’. The Reynolds number describes the ratio of inertial to viscous forces and is defined as $\text{Re}=\frac{\rho UL}{\eta}$, where $\rho$ is the fluid density, $\eta$ its viscosity, and $U$ and $L$ are characteristic velocity and length scales, respectively. In the world of a planktonic cell, $\textrm{Re}\approx 10^{-5}-10^{4}$, thus, in this viscosity-dominated environment, cells overcome and even exploit drag to generate motion Purcell (1977); Lauga and Powers (2009). Cells can achieve this due to the anisotropic drag (defined in Box III) of slender appendages, which allows propulsive forces to be generated perpendicular to the local velocity of the filament Lauga and Powers (2009). Additionally, drag-based thrust relies on the ability of filaments to perform time-periodic deformations that are also time-irreversible motions, crucial for creating a non-zero time-averaged propulsive force Purcell (1977); Lauga and Powers (2009). This can be achieved via the rotation of helical flagella or the whip-like beating of cilia.
@@
Multiscale dynamics. The activity of cellular filamentous appendages encompasses dynamics across multiple scales Gilpin et al. (2020), offering physicists many opportunities to develop frameworks and provide insights into how these dynamics integrate and interact across scales. Nanoscale molecular motors drive shape changes and motions of individual filaments, while these shape changes drive fluid flows Gilpin et al. (2020); Beeby et al. (2020). Resistive force theory (RFT) and slender body theory (SBT) are two hydrodynamic models of fluid-structure interactions for slender bodies within a low Reynolds number environment (see reviews Lauga and Powers (2009); Bondoc-Naumovitz et al. (2023) and references therein). Both are used to estimate the propulsive forces generated by slender appendages based on their waveforms and resulting microswimmer velocities. While RFT is a simpler approximation that neglects long-range hydrodynamic interaction, SBT is more accurate but also more computationally intensive. Mechanical coupling, both via hydrodynamic interactions and the internal cell cytoskeleton, enables coordination of multiple appendages Wan and Poon (2024); Wan (2018), and motivates a range of work that considers cellular appendages as coupled oscillators Uchida et al. (2017); Bruot and Cicuta (2016). Arrays of cilia can synchronise into metachronal waves Poon et al. (2023); Meng et al. (2021); Wan and Poon (2024); Elgeti and Gompper (2013), while the bundling and unbundling dynamics of groups of bacterial flagella are key for their run-and-tumble motility behaviour Grognot and Taute (2021); Wadhwa and Berg (2022). Large numbers of cilia can also group together to form compound ciliary structures, such as the leg-like ‘cirri’ the ciliate Euplotes uses for walking Laeverenz-Schlogelhofer and Wan (2024) and the comb plates of ctenophores that comprise tens of thousands of cilia Jokura et al. (2022). The individual and collective activity of filamentous appendages drive fluid transport, self-propulsion and organism-environment interactions. Groups of microswimmers can display collective behaviours such as bacterial swarming Be’er and Ariel (2019); Wadhwa and Berg (2022) and formation of bioconvection patterns that arise from hydrodynamic instabilities Bees (2020); Schwarzendahl and Mazza (2018); Vachier and Mazza (2019); Schwarzendahl and Mazza (2019, 2022).
@@
Figure 3: Four principles of activation in filamentous cellular appendages. (a) Rotationally actuated filaments are typically rigid helical filaments rotated at their base, such as bacterial flagella seen here in an image reproduced from Turner et al. (2000) of a fluorescently labelled Escherichia coli cell. (b) Linearly actuated bending is the mode of activation found in the cilia of eukaryotes, with typical waveforms consisting of power and recovery strokes. The microalga Chlamydomonas uses two cilia for motility, here shown as three overlaid frames from a movie of its swimming behaviour. (c) Assembly-disassembly activity involves the dynamic extension and retraction of filaments, such as in pili used to generate twitching motility in bacteria and archaea. An example cryo-electron micrograph of a negatively stained wild-type Thermus thermophilus cell shows long flexible pili filaments Neuhaus et al. (2020). (d) Pre-stressed springs respond to changes in intracellular calcium resulting in rapid movements, as seen in the time series of the contractile motion of the stalk of Vorticella Ryu and Matsudaira (2010) and the haptoneme of a haptophyte alga Nomura et al. (2019).
@@
Rotationally actuated filaments. The rotation of passive filaments by cells involves a rotary complex embedded in the cell membrane at the base of the filament where stator and rotary components interact to generate torque that is transmitted to an inert rigid helical filament. There are two main examples of such filaments: the bacterial flagella and the archaea’s archaella Beeby et al. (2020). Their molecular structures differ. The bacterial flagellum is a hollow cylinder self-assembled from the flagellin protein into a a chiral structure Silverman and Simon (1977); Coombs et al. (2002). The archaea’s archaella has a complex structure Cvirkaite-Krupovic et al. (2023). However, they are both $\sim$10-20 nm wide and up to $\sim$10 $\mu$m in length. Archaella use an ATPase to generate torque via ATP hydrolysis Streif et al. (2008); Albers and Jarrell (2015), while the flagella’s rotary complex takes a more structurally intricate form and is powered by the proton motive force Berg (2003); Manson et al. (1977); Wadhwa and Berg (2022). In both cases, the torque generated by the embedded molecular machinery results in a rotation of the connected passive rigid filament that in turn acts as a propellor for cell motility Beeby et al. (2020).
@@
Linearly actuated filaments. In contrast, cilia (sometimes referred to as eukaryotic flagella) are flexible filaments, typically $\sim$250 nm wide and up to $\sim$100 $\mu$m long, and are actuated by linear stepper motors distributed along their entire length Beeby et al. (2020). Cilia have a highly conserved structure across eukaryotes known as the axoneme, which comprises 9 microtubule doublets surrounding a central pair of microtubules Nicastro et al. (2006). Ciliary bending and twisting is powered by dynein motors that walk along the 9 outer doublets. Since the microtubules are usually fixed to the cell body at the base of the cilium, this linear forcing generates differential sliding between neighbouring doublets resulting in bending of the axoneme Lindemann and Lesich (2010); Satir et al. (2014). The mechanism by which dynein motor activity generates ciliary beating is still under debate, with several biophysical models proposed that largely fall under two categories: (1) mechanical feedback within the axoneme regulates dynein activity Lindemann and Lesich (2010); Riedel-Kruse et al. (2007); Cass and Bloomfield-Gadêlha (2023), (2) continuous dynein activity along the flexible filament results in dynamic instabilities Woodhams et al. (2022). Dynein motors are typically distributed along the whole length of the cilium, giving a high degree of flexibility in the types of waveforms that are possible to produce by coordinating the dynein activity in different ways. Compared with the rotation of a rigid filament, which only has a small set of parameters that can be controlled (i.e. speed and direction of rotation), cilia have the potential to display a much more complex and nuanced diversity of waveforms, with calcium playing a central role in the regulation of ciliary activity Eckert (1972). The ability of cilia to change their waveform is critical for enabling ciliated organisms to switch locomotor gaits and perform complex motion Wan and Goldstein (2018); Laeverenz-Schlogelhofer and Wan (2024). Although cilia are highly conserved, there are several examples of variations or additions to the standard 9+2 axoneme structure that exemplify how cilia can be adapted to perform specific functions Ginger et al. (2008). For example, the cilium of the parasite Trypanosoma brucei has a paraflagellar rod, a lattice-like filament that runs along the length of the axoneme, thought to affect axonemal beating by imposing mechanical constraints and contributing to signaling pathways Langousis and Hill (2014). Excavates, species of phagotrophic flagellates, are characterised by possessing vaned cilia within a ventral groove whose confined beating is thought to be particularly efficient for generating feeding currents Suzuki-Tellier et al. (2024).
@@
Assembly-disassembly dynamics. An entirely different principle of activity is that seen in the type IV pili of bacteria and archaea, in which cycles of extension, surface binding and retraction enable cells to perform functions such as twitching motility, surface sensing and DNA uptake Craig et al. (2019). This activity is enabled by ATP hydrolysis, which powers the rapid assembly and disassembly of pilin subunits Ellison et al. (2022).
@@
Pre-stressed springs. Lastly, pre-stressed springs composed of filaments and tubules power many of the fastest cellular movements Mahadevan and Matsudaira (2000). There are two main types: a microtubule-based system found in some microalgae (i.e. the haptoneme in haptophytes) Nomura et al. (2019) and a myoneme based system found in some ciliates (e.g. the spasmoneme of vorticella) Ryu et al. (2017); Floyd et al. (2023). In both cases, rapid coiling is thought to be driven by an increase in Ca2+ concentration, which changes the energetically preferred conformation of the filament from elongated to coiled Kawachi and Inouye (1994); Floyd et al. (2023). For the haptoneme, some evidence suggests that this is driven by conformational changes in microtubule associated proteins that bind with Ca2+ Nomura et al. (2019), while for the spasmoneme it is hypothesised to be a combined electrostatic and entropic effect Mahadevan and Matsudaira (2000).
@@
Elastodynamic properties. A key mechanical parameter for characterising the movements of cellular appendages is the bending stiffness (see Box II), which can be estimated from the deformation induced by forces and torques applied using techniques such as optical tweezers Darnton and Berg (2007), microneedles Xu et al. (2016), magnetic beads Hill et al. (2010) or microfluidic flow devices Shen et al. (2022). Using such approaches the bending stiffness for the bacterial flagella has been estimated to be $1-1000$ pN $\mu$m2 (see table S1 in Shen et al. (2022)), while estimates for eukaryotic motile cilia fall within the 300-6000 pN $\mu$m2 range Okuno and Hiramoto (1979); Hill et al. (2010); Xu et al. (2016); Ishijima and Hiramoto (1994). The Young’s modulus $E$ is generally estimated indirectly from experimentally measured values of the bending stiffness. This requires approximating the second area moment $I$, which is challenging when a filament is not easily defined as a solid or hollow cylinder, such as in the case for cilia. For motile cilia, two values are commonly used: $I=1.3\times 10^{-31}$ m4 Holwill (1965) and $I=3\times 10^{-29}$ m4 Baba (1972), the latter assumed more relevant for tighter connections between the microtubulues of the axoneme Okuno and Hiramoto (1979). This gives a large range of estimates $E\sim(0.01-120)$ GPa.
@@
Appendages as cellular antennae. In addition to their role in motility and fluid transport, appendages can also play important roles in sensory reception and signal transduction that benefit from the geometric and mechanical properties of their filamentous form Marshall and Nonaka (2006); Hickey et al. (2021); R. Ferreira et al. (2019); Craig et al. (2019). Cilia often possess many sensory structures such as chemical receptors and mechanically-sensitive ion channels that enable them to act as the cell’s antenna Marshall and Nonaka (2006); Bloodgood (2010). Filamentous appendages can extend far beyond the cell surface, reaching more well-mixed and therefore more ‘representative’ regions of the surrounding environment, making them effective as chemosensory probes Marshall and Nonaka (2006). Additionally, their high aspect-ratio and high surface area to volume ratio, improves their capture rate and chemosensory effectiveness Hickey et al. (2021). Their elongated shape also means that filamentous appendages can be considered as extended cantilevers, facilitating their sensitivity to mechanical stimuli R. Ferreira et al. (2019).

### v2

@@ III Cellular appendages
In a letter sent to the Royal Society of London in 1676, Antonie van Leeuwenhoek reported the discovery of protozoa and the observation of cilia and flagella. He wrote: “I also discovered a second sort of animalcules, whose figure was an oval, … provided with diverse incredibly thin little feet, or little legs [cilia], which were moved very nimbly …, and wherewith they brought off incredibly quick motions” (cf. translation of letter 18 [228, 229]).
@@
Today, our language has changed a bit, and we describe cilia and flagella as slender, hair-like appendages that extend from the main cell body and perform diverse functions, such as enabling cell movement [230, 231], generating flows [232, 233, 234], sensing environmental stimuli [235, 236] and attachment to surfaces [237, 238]. While there are a myriad of different filamentous appendages found in biology, here our focus is on those most widely studied, with the aim to capture the diversity of their structures, their main properties, mechanisms and functionalities.
@@
Across the tree of life, the most well studied cellular appendages are those used for motility, in particular the flagella in prokaryotes and cilia in eukaryotes. These filaments whip, wave, or rotate through the action of molecular motors that transform the cell’s chemical energy into mechanical energy [230, 10]. The resulting non-reciprocal motion enables net propulsion within the cell’s microscale, viscosity-dominated environment [239, 11].
@@
A low Reynolds number world. At the microscale in which these filaments operate, inertia is insignificant and viscous forces dominate. In fluid dynamics, this is the world of ‘Low Reynolds numbers’. The Reynolds number describes the ratio of inertial to viscous forces and is defined as $\text{Re}=\frac{\rho UL}{\mu}$, where $\rho$ is the fluid density, $\mu$ its viscosity, and $U$ and $L$ are characteristic velocity and length scales, respectively. In the world of planktonic cells and filamentous appendages, typically $\textrm{Re}\sim 10^{-5}-10^{-2}$. In some extreme cases, where motions transiently reach ultrafast velocities, the local Re can be much higher, up to $\sim 10^{3}$ [240]. Cellular scale motions are thus mostly within a low Re regime. In this viscosity-dominated environment, cells overcome and even exploit drag to generate motion [239, 11]. Cells can achieve this due to the anisotropic drag (defined in Box III) of slender appendages, which allows propulsive forces to be generated perpendicular to the local velocity of the filament [11]. Additionally, drag-based thrust relies on the ability of filaments to perform time-periodic deformations that are also non-reciprocal, that is, the sequences of motion differ under a time-reversal [239, 11], crucial for creating a non-zero time-averaged propulsive force [239, 11]. This can be achieved via the rotation of helical flagella or the whip-like beating of cilia.
@@
Multiscale dynamics. The activity of cellular filamentous appendages encompasses dynamics across multiple scales [10], offering physicists many opportunities to develop frameworks and provide insights into how these dynamics integrate and interact across scales. Nanoscale molecular motors drive shape changes and motions of individual filaments, while these shape changes drive fluid flows [10, 230]. Resistive force theory (RFT) and slender body theory (SBT) are two hydrodynamic models of fluid-structure interactions for slender bodies within a low Reynolds number environment (see reviews [11, 231] and references therein). Both are used to estimate the propulsive forces generated by slender appendages based on their waveforms and resulting microswimmer velocities. While RFT is a simpler approximation that neglects long-range hydrodynamic interaction, SBT is more accurate but also more computationally intensive. The basic idea of SBT for Stokes flow is to replace the disturbance due to a body with a suitably chosen line distribution of Stokeslets [241]. Because the Stokes equation $\mu\nabla^{2}\bm{u}-\nabla p=0$, $\nabla\cdot\bm{u}=0$ is linear, a Green’s function exists; for a point-like forcing $\bm{F}\delta(\bm{r})$ at the origin in an unbounded fluid, the fluid velocity at point $\bm{r}$ is
${u}_{i}(\bm{r})=\frac{F_{j}}{8\pi\mu}\left(\frac{\delta_{ij}}{|\bm{r}|}+\frac{r_{i}r_{j}}{|\bm{r}|^{3}}\right)\,,$ (17)
where $\delta_{ij}$ is the Kronecker delta, and Einstein’s summation convention is used. The expression in Eq.(17) is called a ‘Stokeslet’. In the presence of a body with length $L$ much larger than its diameter $2a$, the key idea of SBT is to approximate the flow with a line density of Stokeslets
${u}_{i}^{\mathrm{SBT}}(\bm{r})=$ $\int_{0}^{L}\frac{{F}_{j}(s)}{8\pi\mu}\Bigg{[}\frac{\delta_{ij}}{|\bm{r}-\bm{R}(s)|}$ (18)
$+\frac{({r}_{i}-{R}_{i}(s))({r}_{j}-{R}_{j}(s))}{|\bm{r}-\bm{R}(s)|^{3}}\Bigg{]}ds\,,$ (19)
where $s$ and and $\bm{R}(s)$ are the arc length and position vector along the filament, respectively [241, 242, 11]. SBT has been successufully used to reproduce experimental results with better precision than RFT [243].
Mechanical coupling, both via hydrodynamic interactions and the internal cell cytoskeleton, enables coordination of multiple appendages [244, 245], and motivates a range of work that considers cellular appendages as coupled oscillators [246, 247]. Arrays of cilia can synchronise into metachronal waves [248, 249, 244, 250], while the bundling and unbundling dynamics of groups of bacterial flagella are key for their run-and-tumble motility behaviour [251, 252]. Large numbers of cilia can also group together to form compound ciliary structures, such as the leg-like ‘cirri’ the ciliate Euplotes uses for walking [253] and the comb plates of ctenophores that comprise tens of thousands of cilia [254]. The individual and collective activity of filamentous appendages drive fluid transport, self-propulsion and organism-environment interactions. Groups of microswimmers can display collective behaviours such as bacterial swarming [255, 252] and formation of bioconvection patterns that arise from hydrodynamic instabilities [256, 257, 258, 259, 260].
@@
Figure 3: Four principles of activation in filamentous cellular appendages. (a) Rotationally actuated filaments are helical filaments rotated at their base, such as bacterial flagella seen here in an image reproduced from [261] of a fluorescently labelled Escherichia coli cell. (b) Linearly actuated bending is the mode of activation found in the cilia of eukaryotes, with typical waveforms consisting of power and recovery strokes. The microalga Chlamydomonas uses two cilia for motility, here shown as three overlaid frames from a movie of its swimming behaviour. (c) Assembly-disassembly activity involves the dynamic extension and retraction of filaments, such as in pili used to generate twitching motility in bacteria and archaea. An example cryo-electron micrograph of a negatively stained wild-type Thermus thermophilus cell shows long flexible pili filaments [262]. (d) Pre-stressed springs respond to changes in intracellular calcium resulting in rapid movements, as seen in the time series of the contractile motion of the stalk of Vorticella [237] and the haptoneme of a haptophyte alga [263].
@@
$\bullet$ Rotationally actuated filaments. The rotation of passive filaments by cells involves a rotary complex embedded in the cell membrane at the base of the filament where stator and rotary components interact to generate torque that is transmitted to a helical filament [264]. There are two main examples of such filaments: the bacterial flagella and the archaea’s archaella [230]. Their molecular structures differ. The bacterial flagellum is a hollow cylinder self-assembled from the flagellin protein into a chiral structure and connected to the rotary motor via a short flexible hook at its base [265, 266]. The archaea’s archaella have a solid multi-stranded helical structure [267]. However, they are both $\sim$10-20 nm wide and up to $\sim$10 $\mu$m in length. Archaella use an ATPase to generate torque via ATP hydrolysis [268, 269], while the flagella’s rotary complex takes a more structurally intricate form and is powered by the proton motive force [270, 271, 252]. In both cases, the torque generated by the embedded molecular machinery results in a rotation of the connected helical filament that in turn acts as a propeller for cell motility [230].
Although these helical filaments are often considered rigid, flexibility plays a crucial functional role [272]. Flexibility of the basal hook enables flagella bundling [273] and, through a buckling instability, allows abrupt directional changes via an off-axis motion of the flagellum known as a ‘flick’ [274]. Intuitively, peritrichously-flagellated bacteria should have negligible swimming speeds because of the flagella distribution around the cell body. However, an elastohydrodynamic bending instability is responsible for a conformational change in the hook bending and results in a net propulsive force [275, 276]. Additionally, flagella can undergo polymorphic transformations —shape changes driven by cooperative conformational changes in flagellin— depending on factors like torsional load, pH, temperature and ionic strength [277]. These expand the functional versatility of flagella, such as enabling run-and-tumble behaviour and escape from traps [261, 278].
@@
$\bullet$ Linearly actuated filaments. In contrast, cilia (sometimes referred to as eukaryotic flagella) are typically $\sim$250 nm wide and up to $\sim$100 $\mu$m long, and are actuated by linear stepper motors distributed along their entire length [230]. Cilia have a highly conserved structure across eukaryotes known as the axoneme, which comprises 9 microtubule doublets surrounding a central pair of microtubules [279]. Ciliary bending and twisting is powered by dynein motors that walk along the 9 outer doublets. Since the microtubules are usually fixed to the cell body at the base of the cilium, this linear forcing generates differential sliding between neighbouring doublets resulting in bending of the axoneme [280, 281]. The mechanism by which dynein motor activity generates ciliary beating is still under debate, with several biophysical models proposed that largely fall under two categories: (1) mechanical feedback within the axoneme regulates dynein activity [280, 282, 283], (2) continuous dynein activity along the cilium results in dynamic instabilities [284]. Dynein motors are typically distributed along the whole length of the cilium, giving a high degree of flexibility in the types of waveforms that are possible to produce by coordinating the dynein activity in different ways. Compared with the basally actuated rotation of a helical filament, which only has a small set of parameters that can be controlled (e.g. speed and direction of rotation), cilia have the potential to display a much more complex and nuanced diversity of waveforms, with calcium playing a central role in the regulation of ciliary activity [285]. The ability of cilia to change their waveform is critical for enabling ciliated organisms to switch locomotor gaits and perform complex motion [286, 253]. Although cilia are highly conserved, there are several examples of variations or additions to the standard 9+2 axoneme structure that exemplify how cilia can be adapted to perform specific functions [287]. For example, the cilium of the parasite Trypanosoma brucei has a paraflagellar rod, a lattice-like filament that runs along the length of the axoneme, thought to affect axonemal beating by imposing mechanical constraints and contributing to signaling pathways [288]. Excavates, species of phagotrophic flagellates, are characterised by possessing vaned cilia within a ventral groove whose confined beating is thought to be particularly efficient for generating feeding currents [289].
@@
$\bullet$ Assembly-disassembly dynamics. An entirely different principle of activity is that seen in the type IV pili of bacteria and archaea, in which cycles of extension, surface binding and retraction enable cells to perform functions such as twitching motility, surface sensing and DNA uptake [290]. This activity is enabled by ATP hydrolysis, which powers the rapid assembly and disassembly of pilin subunits [291].
@@
$\bullet$ Pre-stressed springs. Lastly, pre-stressed springs composed of filaments and tubules power many of the fastest cellular movements [292]. There are two main types: a microtubule-based system found in some microalgae (i.e. the haptoneme in haptophytes) [263] and a myoneme based system found in some ciliates (e.g. the spasmoneme of vorticella) [293, 294]. In both cases, rapid coiling is thought to be driven by an increase in Ca2+ concentration, which changes the energetically preferred conformation of the filament from elongated to coiled [295, 294]. For the haptoneme, some evidence suggests that this is driven by conformational changes in microtubule associated proteins that bind with Ca2+ [263], while for the spasmoneme it is hypothesised to be a combined electrostatic and entropic effect [292].
@@
Elastodynamic properties. A key mechanical parameter for characterising the movements of cellular appendages is the bending stiffness (see Box II), which can be estimated from the deformation induced by forces and torques applied using techniques such as optical tweezers [296], microneedles [297], magnetic beads [298] or microfluidic flow devices [82]. Using such approaches the bending stiffness for the bacterial flagella has been estimated to be in the range $1-1000$ pN $\mu$m2 (see Table S1 in [82]), while estimates for eukaryotic motile cilia fall within the 300-6000 pN $\mu$m2 range [85, 298, 297, 86]. The Young’s modulus $E$ is generally estimated indirectly from experimentally measured values of the bending stiffness. This requires approximating the second area moment $I$, which is challenging when a filament is not easily defined as a solid or hollow cylinder, such as in the case for cilia. For motile cilia, two values are commonly used: $I=1.3\times 10^{-31}$ m4 [299] and $I=3\times 10^{-29}$ m4 [300], the latter assumed more relevant for tighter connections between the microtubules of the axoneme [85]. This gives a large range of estimates $E\sim(0.01-120)$ GPa.
@@
Appendages as cellular antennae. In addition to their role in motility and fluid transport, appendages can also play important roles in sensory reception and signal transduction that benefit from the geometric and mechanical properties of their filamentous form [235, 301, 236, 290]. Both motile and non-motile cilia can provide sensory functions, possessing chemical receptors and mechanically-sensitive ion channels that enable them to act as the cell’s antenna [235, 302]. Primary cilia, typically non-motile and structurally distinct from motile cilia, are specialised sensory organelles critical for coordinating signal transduction pathways during development and homeostasis [303]. Filamentous appendages can extend far beyond the cell surface, reaching more well-mixed and therefore more ‘representative’ regions of the surrounding environment, making them effective as chemosensory probes [235]. Additionally, their high aspect-ratio and high surface area to volume ratio, improves their capture rate and chemosensory effectiveness [301]. Their elongated shape also means that filamentous appendages can be considered as extended cantilevers, facilitating their sensitivity to mechanical stimuli [236]. Beyond mechanical interactions, recent work suggests a potential role of primary cilia in the regulation of neuronal and glial activity [304, 305].